OpenStudy (ricardo_guzman129489):

Jina made $357 for 17 hours of work.At the same rate, how much would she make for 8 hours of work?

Mehek (mehek14):

Find per hour 357/7 = ?

OpenStudy (purplefirebulovescali4life):

357 /7 =51

Directrix (directrix):

$357 is to 17 work hours as x dollars is to 8 work hours 357/17 = x/8 x = $168 for 8 hours of work @ricardo_guzman129489

OpenStudy (ricardo_guzman129489):

Thank you so much @Directrix

Directrix (directrix):

You are welcome.